# https://fudan-zvg.github.io/Efficient4D

### v2

@@ Abstract
Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data. Extending image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling, existing methods tend to be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly train a novel 4D Gaussian splatting model with explicit point cloud geometry, enabling real-time rendering under continuous camera trajectories. Extensive experiments on synthetic and real videos show that Efficient4D offers a remarkable 20-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis. For example, Efficient4D takes only 6 mins to model a dynamic object, vs 120 mins by Consistent4D.
@@
4D Generation Gaussian Splatting Video Efficiency
@@ 1 Introduction
Refer to caption  Refer to caption Refer to caption
 Refer to caption  Refer to caption Refer to caption
Input Consistent4D [13] Efficient4D(Ours)
Figure 1: Examples of video-to-4D generation. Input: A brief video of a dynamic object, as represented by 3 frames per case; Output: Generated novel views at different timestamps.
@@
Directly generating 4D object/scene content from text description has been recently attempted [38]. To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation [26] via score distillation sampling [29] and separates the modeling of static scene and its dynamics. Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content. In quest of the aforementioned human’s capability, a recent work [13] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation. However, similar as [38] this method is also slow to train (e.g., 150 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
@@
To address identified limitations, we formulate an efficient video-to-4D two staged object generation method called Efficient4D. In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data. This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer [22]) in tandem with frame interpolation. In the second stage, we use these training data to optimize a novel 4D Gaussian splatting model. This is an extension of the 3D Gaussian splatting [15], originally designed for static 3D scene representation, with the temporal dimension introduced additionally. Using Gaussian representation brings about further computational efficiency gain along with the explicit geometry property, when compared with NeRF based designs (Fig. 1). To tackle the challenging discontinuity between generated frames, we design an inconsistency-aware loss function.
@@
Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem. (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches. (iii) Instead of using NeRF as the 3D representation model, we introduce a novel 4D Gaussian splatting model with the real-time rendering advantage. (iv) Extensive experiments on both synthetic and real videos validate the significant efficiency advantage (e.g., 20$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the quality of novel view synthesis. Also, our method can work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
@@ 2 Related work
3D generation takes two main settigns: text-to-3D and image-to-3D. The pioneering work, DreamFusion [29], introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models. SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [18, 25, 34, 36, 4, 16, 41, 11, 47, 43, 52] and image-to-3D tasks [48, 40, 24, 31] across various dimensions. However, SDS-based approaches often suffer from extended optimization times. Conversely, some efficient methods [27, 14, 21, 20, 22] have emerged. Notably, Point-E [27] and Shap-E [14] train models to directly generate 3D point clouds or meshes. Zero123 [21] focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or [20]. Importantly, SyncDreamer [22] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
@@
Our work is perpendicular to all the above works, where any of them can be deployed in our reconstruction phase. But considering both optimization efficiency and expressive capability, we propose to represent dynamic 3D assets by a set of native 4D scene primitives.
4D generation
@@
There are a few recent works dedicated for more challenging 4D object generation. For instance, MAV3D [38] deals with a text-to-4D problem by training Hexplane [3] with a video diffusion model and SDS loss. Instead of text input, Consistent4D [13] conditions the generation of 4D object over time on a monocular video with richer and more specific information. However, it is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design. To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the training speed by 20$\times$ whilst maintaining the quality of novel view synthesis.
3 Method
Figure 2: Overview of our Efficient4D approach. Given as the input a brief video depicting a dynamic object from a single perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time. Efficient4D comprises two components: (A) Image sequence synthesis across views and timestamps, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images. (B) A novel 4D Gaussian representation model (d) that represents the scene with a number of Gaussian points. It can be trained efficiently and robustly under the confidence-aware (e) supervision on the generated image matrix.
Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video. As illustrated in Fig. 2, it comprises two key components:
  • An image synthesis pipeline (Sec. 3.1) generates images across views and timestamps, ensuring sufficient geometry and temporal consistency.
  • A novel 4D Gaussian representation model (Sec. 3.2) efficiently utilizes the synthetic images for accurate dynamic object reconstruction and novel view synthesis.
3.1 Image synthesis across views and timestamps
Due to the difficulty of obtaining calibrated 4D scans, our approach involves the direct generation of high-quality consistent 4D data from a single-view video which is much easier to capture (Fig. 2(A)). Specifically, we seek to produce a $M\times N$ image matrix $\mathcal{D}=\{e_{ij}\}_{i,j=1}^{M,N}$ representing 2D images with geometrical and temporal consistency. Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Fig. 2(c)). This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation [37, 22], offering comprehensive information for modeling a dynamic object.
To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video (Fig. 2(a)) and proceed to generate the remaining views. Our task involves generating multi-view consistent images from a single image for each row. Existing image-to-3D methods, such as SyncDreamer [22], can be leveraged for this purpose. However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images. To address this issue, we propose an enhanced version of SyncDreamer with improved temporal continuity, referred to as SyncDreamer-T.
To impose temporal consistency, the information from $\mathcal{V}$ at different timestamps is aggregated using a time-synchronous spatial volume we design here (Fig. 2(b)). A smoothing filter layer is introduced into the spatial volumes of different frames/timestamps, incorporating a weight vector $\mathbf{w}=(w_{-k},\dots,w_{0},\dots,w_{k})$. At each denoising step, time-synchronized spatial volumes $\tilde{\mathcal{V}}_{i}$ for each input frame $i\in\{1,2,\dots,M\}$ are constructed as:
$\tilde{\mathcal{V}}_{i}=\sum_{j=-k}^{k}w_{i+j}\mathcal{V}_{i+j}.$ (1)
This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, enhancing temporal consistency. With this time-synchronized spatial volumes, the proposed SyncDreamer-T is entirely training-free established on the pretrained SyncDreamer.
To further improve temporal resolution, video frame interpolation (e.g., RIFE [12]) is applied after generating the image matrix $\mathcal{D}$ as described above. The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames. This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
Refer to caption Refer to caption
Input Feature fusion
Analysis on temporal synchronization
For a clearer insight into our temporal synchronization design, we undertake a simplified experiment involving two input frames, each with its feature volumes labeled as $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. The fusion process is carried out by combining them as follows:
$\tilde{\mathcal{V}}_{i}=(1-w)\mathcal{V}_{i}+w\mathcal{V}_{3-i},$ (2)
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion" as illustrated in Fig. 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches 0.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio between 0.25 and 0.4, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry. Note that the results in Fig. 3 are not sensitive to the choice in this range.
3.2 4D generation through reconstruction
Aiming for 3D dynamic object modeling, discrete images not suffice. Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$. For efficient modeling, we formulate a novel 4D Gaussian representation model (Fig. 2(B)), departing from previous slow-to-train 4D reconstruction models [8].
4D Gaussian representation model
@@ 3.1 Image synthesis across views and timestamps
Our model builds upon the 3D Gaussian splatting technique introduced in [15], originally designed for static scene representation. We extend it to address the complexities of dynamic scenes, enabling the real-time synthesis of high-fidelity views in 4D space. The model represents each Gaussian defined as:
@@
In contrast to the 3D counterpart, our formulation incorporates the time dimension, allowing coherent modeling of dynamics in both space and time.
Rendering
@@
Optimization
@@
In training, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian. The original objective function, as presented in [15], assumes clean training data, which may not less valid for synthetic data with inherent imperfections. To address this, we introduce a confidence-aware objective loss formulation defined as:
$\mathcal{L}=\lambda_{\text{rgb}}\mathcal{C}_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{ssim}}\mathcal{C}_{\text{ssim}}\mathcal{L}_{\text{ssim}},$ (9)
@@
where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss [44], $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter, and $\mathcal{C}_{\text{rgb/ssim}}$ is the confidence score of a generated image $\mathbf{I}$ calculated as
$\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}})$ (10)
@@
where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by optical flow. In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions.

### v3

# Efficient4D: Fast Dynamic 3D Object Generation from a Single-view Video

@@ Abstract
Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data. An intuitive approach is to extend previous image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling. However, this approach would be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly reconstruct the 4D content through a 4D Gaussian splatting model. Importantly, our method can achieve real-time rendering under continuous camera trajectories. To enable robust reconstruction under sparse views, we introduce inconsistency-aware confidence-weighted loss design, along with a lightly weighted score distillation loss. Extensive experiments on both synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis. For example, Efficient4D takes only 10 minutes to model a dynamic object, vs 120 minutes by the previous art model Consistent4D. Our code is publicly available at https://github.com/fudan-zvg/Efficient4D.
@@
4D Generation, Efficiency, Diffusion Model, Gaussian Splatting, Video
@@ 1 Introduction
Directly generating 4D object/scene content from text description has been recently attempted [13]. To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation [10] via score distillation sampling [1] and separates the modeling of static scene and its dynamics. Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content. In quest of the aforementioned human’s capability, a recent work [14] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation. However, similar as [13] this method is also slow to train (e.g., 120 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
@@
To address identified limitations, we formulate an efficient video-to-4D two staged object generation method called Efficient4D. In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data. This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer [8]) in tandem with frame interpolation. In the second stage, we use these training data to optimize a 4D Gaussian splatting model [15]. This is an extension of the 3D Gaussian splatting [16], originally designed for static 3D scene representation, with the temporal dimension introduced additionally, allowing for real-time rendering under continuous camera trajectories. Using Gaussian representation brings about further computational efficiency gain, when compared with NeRF based designs (Figure 1). To tackle the challenging discontinuity between the generated sparse frames, we design an inconsistency-aware loss function based on whether there is confidence in the consistency of a frame with its adjacent frames. Along with a lightly weighted score distillation sampling loss for smooth viewpoint transitions, this technique enables robust 4D reconstruction. Notably, although our method takes a single-view video as input, Efficient4Dcan be easily extended to image-to-4D task by leveraging a image-to-video diffusion model [17].
@@
Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem. (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches. We also extends Efficient4D to image-to-4D task. (iii) We introduce a inconsistency-aware confidence-weighted loss for reconstructing a 4D Gaussian splatting model using the generated training data efficiently and robustly. (iv) Extensive experiments on both synthetic and real data validate the significant efficiency advantage (e.g., 10$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the quality of novel view synthesis. Also, our method can work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
Input Consistent4D [14]: 120 mins Efficient4D (Ours): 10 mins
Input 4DGen [18]: 130 mins Efficient4D (Ours): 10 mins
Input STAG4D [19]: 70 mins Efficient4D (Ours): 10 mins
Figure 1: Examples of video-to-4D generation. Input: A brief video of a dynamic object, as represented by 2 frames per case; Output: Generated novel views at different timestamps. The generation time is also shown for each method. More visualized comparisons can be found in Figure 5.
@@ 2 Related work
3D generation takes two main settigns: text-to-3D and image-to-3D. The pioneering work, DreamFusion [1], introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models. SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [2, 20, 21, 22, 3, 23, 24, 25, 26, 4, 27] and image-to-3D tasks [28, 29, 30, 31] across various dimensions. However, SDS-based approaches often suffer from difficulty to converge and extended optimization times. Conversely, some efficient methods [32, 33, 7, 34, 8] have emerged. Notably, Point-E [32] and Shap-E [33] train models to directly generate 3D point clouds or meshes. Zero123 [7] focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or [34]. Importantly, SyncDreamer [8] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
[…]
@@
Our work is perpendicular to all the above works, where any of them can be deployed in our reconstruction phase. But considering both optimization efficiency and expressive capability, we choose to represent dynamic 3D assets by a set of native 4D scene primitives, which is proved to be superior in 4DGS [15].
[…]
@@
4D generation There are a few recent works dedicated for more challenging 4D object generation. For instance, MAV3D [13] deals with a text-to-4D problem by training Hexplane [41] with a video diffusion model and SDS loss. Instead of text input, Consistent4D [14] conditions the generation of 4D object over time on a monocular video with richer and more specific information. However, it is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design. To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the training speed by 20$\times$ whilst maintaining the quality of novel view synthesis.
3 Preliminary
[…]
4D Gaussian splatting
@@ 3 Preliminary
The 4D Gaussian splatting (4DGS [15]) builds upon the 3D Gaussian splatting technique introduced in [16], originally designed for static scene representation. To address the complexities of dynamic scenes, 4DGS represents each Gaussian defined as:
@@
[…]
Score distillation sampling Score distillation sampling (SDS) is first introduced by DreamFusion [1], which is used to distill the knowledge from a pretrained diffusion model $\epsilon_{\phi}$. Specifically, given an image $I$ rendered from a scene representation (e.g. 3DGS) parameterized by $\theta$, the gradient of SDS loss is calculated as:
$\nabla_{\theta}\mathcal{L}_{\mathrm{SDS}}(\phi,I_{t})=\mathbb{E}\left[w(t)\left(\epsilon_{\phi}(I_{t};t,c)-\epsilon\right)\frac{\partial I_{t}}{\partial\theta}\right],$ (7)
where $I_{t}$ is the perturbed image with noise $\epsilon$ at time step $t$, and $c$ is the condition (e.g. one frame of the input video in this paper).
4 Method
Figure 2: Overview of our Efficient4D approach. Given as the input (a) a brief video depicting a dynamic object from a single perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time. Efficient4D comprises two components: (A) Image sequence synthesis through (b) time-synchronous spatial volumes, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images. (B) 4D Reconstruction using the generated images in (A). The 4D Gaussian representation can be trained efficiently and robustly under the confidence-weighted loss $\mathcal{L}_{\text{img-conf}}$ and the low-weighted SDS loss $\mathcal{L}_{SDS}$.
Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video. The input single-view video can be either provided by user or generated by a video generation model. The latter approach extends the application of our method beyond video-to-4D. For example, we can also achieve image-to-4D through a image-to-video diffusion model [17]. As illustrated in Figure 2, it comprises two key components:
  • An image synthesis pipeline (Figure 2(A)) generates images across views and timestamps, ensuring sufficient geometry and temporal consistency.
  • An efficient and robust reconstruction process (Figure 2(B)) efficiently utilizes the synthetic images for accurate dynamic object reconstruction and novel view synthesis.
We will elaborate on these components in Section 4.1 and 4.2, respectively.
4.1 Image synthesis across views and timestamps
Due to the difficulty of obtaining calibrated 4D scans, our approach involves the direct generation of high-quality consistent 4D data from a single-view video which is much easier to capture (Figure 2(a)). Specifically, we seek to produce a $M\times N$ image matrix $\mathcal{D}=\{e_{ij}\}_{i,j=1}^{M,N}$ representing 2D images with geometrical and temporal consistency. Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Figure 2(c)). This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation [47, 8], offering comprehensive information for modeling a dynamic object.
To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video and proceed to generate the remaining views. Our task involves generating multi-view consistent images from a single image for each row. Existing image-to-3D methods, such as SyncDreamer [8], can be leveraged for this purpose. However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images. To address this issue, we propose an enhanced version of SyncDreamer with improved temporal continuity, referred to as SyncDreamer-T.
To impose temporal consistency, the information from $\mathcal{V}$ at different timestamps is aggregated using a time-synchronous spatial volume we design here (Figure 2(b)). A smoothing filter layer is introduced into the spatial volumes of different frames/timestamps, incorporating a weight vector $\mathbf{w}=(w_{-k},\dots,w_{0},\dots,w_{k})$ which serves as the smooth filter. At each denoising step, time-synchronized spatial volumes $\tilde{\mathcal{V}}_{i}$ for each input frame $i\in\{1,2,\dots,M\}$ are constructed as:
$\tilde{\mathcal{V}}_{i}=\sum_{j=-k}^{k}w_{i+j}\mathcal{V}_{i+j}.$ (8)
This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, thus enhancing temporal consistency. With this time-synchronized spatial volumes, the proposed SyncDreamer-T is entirely training-free established on the pretrained SyncDreamer.
To further improve temporal resolution, video frame interpolation (e.g., RIFE [48]) can be applied after generating the image matrix $\mathcal{D}$. The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames. This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
Refer to caption Refer to caption
Input Feature fusion
[…]
Analysis on temporal synchronization For a clearer insight into our temporal synchronization design, we undertake a simplified experiment involving two input frames, each with its feature volumes labeled as $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. The fusion process is carried out by combining them as follows:
$\tilde{\mathcal{V}}_{i}=(1-w)\mathcal{V}_{i}+w\mathcal{V}_{3-i},$ (9)
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion" as illustrated in Figure 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches $0.5$, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio just above $0.5$, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry. Also note that the results in Figure 3 are not sensitive to the choice.
4.2 4D generation through reconstruction
Aiming for 3D dynamic object modeling, discrete images not suffice. Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$. For efficient modeling, we have formulated the 4DGS representation model in Section 3, departing from previous slow-to-train 4D reconstruction models [40].
[…]
@@ 4.1 Image synthesis across views and timestamps
Optimization In the training of 4DGS, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian. The original loss function, as presented in [15], involves both RGB loss and SSIM loss with balancing weights fixed. Specifically, the loss formulation is defined as:
$\mathcal{L}_{\text{img}}=\lambda_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{ssim}}\mathcal{L}_{\text{ssim}},$ (10)
@@
where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss [49], and $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter.
However, such optimization approach assumes clean training data, which may not valid for synthetic data with inherent imperfections. To address this, we first introduce a inconsistency-aware loss formulation with adaptive balancing weights:
$\lambda_{\text{rgb}}=\tilde{\lambda}_{\text{rgb}}\mathcal{C}_{\text{rgb}},\quad\lambda_{\text{ssim}}=\tilde{\lambda}_{\text{ssim}}\mathcal{C}_{\text{ssim}},$ (11)
where $\tilde{\lambda}_{\text{rgb/ssim}}$ is fixed weight and $\mathcal{C}_{\text{rgb/ssim}}$ is the adaptive confidence score of a generated image $\mathbf{I}$ calculated as
$\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}}),$ (12)
@@
where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by optical flow. In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions, thus enhancing overall reconstruction quality.
The confidence design can guarantee temporal consistency, but achieving high quality of novel view synthesis still remains challenging due to the sparsity of the generated images. Therefore, we also incorporate SDS loss with a relatively small weight $\lambda_{\text{SDS}}$ for smooth transition across different supervised views. We use the image-to-3D diffusion model [17] in SDS loss conditioned on the frame of input video at each timestamp. The total loss function is expressed by:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{img-conf}}+\lambda_{\text{SDS}}\mathcal{L}_{\text{SDS}}$ (13)
with $\mathcal{L}_{\text{img-conf}}=\tilde{\lambda}_{\text{rgb}}\mathcal{C}_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\tilde{\lambda}_{\text{ssim}}\mathcal{C}_{\text{ssim}}\mathcal{L}_{\text{ssim}}$.